OpenStudy (anonymous):

Factor the GCF from each term in the expression. 7b(a-9)-4(a-9)..7b(a-9)-4(a-9)=??

OpenStudy (anonymous):

What is common?

OpenStudy (anonymous):

I will say (a-9). Why?

OpenStudy (anonymous):

because that is in all of them right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so what would be my next step?

OpenStudy (anonymous):

7b(a-9)-4(a-9)..7b(a-9)-4(a-9)= (a-9) (?)

OpenStudy (anonymous):

ok im lost now

OpenStudy (anonymous):

7b(a-9)-4(a-9)..7b(a-9)-4(a-9)= (a-9) (7b -4 -7b -4) What is ..7b(a-9)? is it -7b(a-9)?

OpenStudy (anonymous):

yes so that would be my answer. -7b(a-9)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

7b(a-9)-4(a-9)..7b(a-9)-4(a-9)= (a-9) (7b -4 -7b -4) 7b(a-9)-4(a-9)..7b(a-9)-4(a-9)= (a-9) (-8)=8(9-a)

OpenStudy (anonymous):

so 8(9-a) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the final answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

or would i multiply 8 * 9

OpenStudy (anonymous):

No. you should leave it in factored form.

OpenStudy (anonymous):

i eneterd the 8(9-a) and it was wrong it said

OpenStudy (anonymous):

That is the answer to the problem that you wrote down.

OpenStudy (anonymous):

you realize the problem is just to the two .. right

OpenStudy (anonymous):

it's not both i just typed it twice sorry

OpenStudy (anonymous):

so the problem is 7b(a-9)-4(a-9)

OpenStudy (anonymous):

The answer is (7b-4)(a-9)